OVERTONE VIBRATIONAL SPECTROSCOPY AND DYNAMICS IN H_{2}-H_{2}O COMPLEXES: A COM-BINED THEORETICAL AND EXPERIMENTAL STUDY

\underline{MICHAEL P. ZIEMKIEWICZ}, CHRISTIAN PLUETZER and DAVID J. NESBITT , JILA, University of Colorado and National Institute of Standards and Technology, Boulder, Colorado, USA; YOHANN SCRIBANO, CNRS-Universite de Bourgogne, Dijon, France; ALEXANDRE FAURE, CNRS, Institut de Planetologie et dAstrophysique de Grenoble(IPAG), France; and AD VAN DER AVOIRD, Radboud University, Nijmegen, The Netherlands.

We present spectroscopic studies on the H_{2}-H_{2}O complex, providing information on the binding of these astrochemicallysignificant molecules. Understanding this interaction is crucial for further understanding the production of H_{2} from atomic hydrogen on interstellar dust grain surfaces, which are often covered with an icy mantel. We generate these clusters using a slit supersonic jet expansion which cools to below 5 K, a temperature which is sufficiently low for observing this species whose binding energy is 60 cm^{−1}. The experimental detection method consists of an infrared vibrational excitation followed by a 193 nm pulse to photolyze H_{2}O and laser induced fluorescence on the OH product. In conjunction with theoretical studies presented in the accompanying talk, we identify two overtone bands associated with excitation of the |02−⟩ ← |00+⟩ transition in the water moiety. Theory and experiment agree to within 0.2 cm^{−1}, which attests to the accuracy of the method used to make predictions, as shown in the accompanying talk. Additional time and frequency domain measurements are performed with this three laser experiment to obtain information on predissociation dynamics, where internal H_{2}O vibration is transferred to the intermolecular coordinate. Theoretically-generated excited state wavefunctions are used to obtain further insight into the significant dependence of predissociation lifetime on the upper state in the infrared cluster transition.